# Electron transport in disordered insulating lattice under nonlinear electric field

### v2

@@ III Results
III.1 Electron propagation in Disordered Lattice
@@
We compute $C(x)$ in the finite-lattice calculation numerically by computing the disorder-averaged correlation of the non-diagonal matrix elements terms of the retarded Green’s function at the local Fermi energy $-mE$ averaged over each $m$-th site, since the transport is carried over states close to the Fermi energy. Here, $x$ denotes the spatial index ($0\leq x\leq 150$) across the 1D lattice and $M=201$ denotes the number of sites chosen for spatial averaging around the center of the lattice. This quantity $C(x)$ is averaged over multiple disorder configurations ($\sim 1000$ different configurations).
For strongly localized states, this quantity decays exponentially and for a delocalized or extended state, it spreads over the entire lattice. Fig. 2(a) depicts the behavior of $C(x)$ for increasing electric field at the disorder strength $W=0.4$ inside the VRH regime. At low electric fields, the wave functions are typically localized; but at higher fields, the wave function shows evidence for the Bloch oscillations. Although the Bloch oscillation localizes the electron motion via the Brillouin zone averaging, it is only possible when electron traverses the Brillouin zone. Therefore, the emergence of the Bloch oscillation can be taken as a sign of the delocalization. The spread of $C(x)$ behaves inversely proportional to the electric field and at high fields $E\gtrsim 0.4$, $C(x)$ extends to the distance $x_{B}=4t/E$ (as marked by an arrow for $E=0.07$). The width of the Bloch oscillation can be easily understood as the travel distance of an electron driven by an electric field as
@@
with the group velocity $v_{g}(k)=2t\sin(k)$ and the semi-classical equation of motion $\hbar\dot{k}=eEa$. This behavior corroborates with the discussion [35] that, at high fields, the agreement of the CPA and the finite-lattice spectral functions indicates the tendency of delocalization.
@@
Figure 3: (a) Conductivity vs electric field ($\Gamma=0.0005$, $N=501$, $t=1$, $T=0.01$). The average conductivity slowly increases with the electric field. In the low electric field region the conductivity strongly varies with the disorder strengths [$W\in\{0.2,0.3,0.35,0.4,0.45,0.5,0.6,0.7,0.8,1.0,1.2\}$: increasing in order indicated by the arrow in 3(b)] with a rapid rise in the VRH regime $0.50.3<W<$ 0.5. The black solid and the dashed line show the numerical and analytic behaviors [see Eq. (32)], respectively, at zero disorder. (b) Mobility vs disorder strength ($\Gamma=0.0005$, $N=501$, $t=1$, $T_{B}=0.01$) for different electric fields. The average mobility at low electric fields $E<0.04$ sharply rises in the disorder range $\Delta<W\lesssim 0.5$ marked by the black dashed lines for the VRH regime. Beyond this disorder range ($W>0.5$) the mobility typically follows roughly the Fermi golden-rule ($\sim W^{-2}$) shown by the black dashed line. (c) CPA calculation of conductivity. While the overall behavior is consistent with the finite-lattice calculation, the high-disorder limit strongly overestimates the conductivity. (d) CPA calculation of mobility. Despite the similarity of the overall lineshape, the CPA overestimates the mobility and fails to produce the correct trends in the high-$W$ limit.
@@
[…]
We make a note here that, the drift distance is the travel distance of wave-propagation under the effect of an external field during the charge carrier’s lifetime. This is conceptually distinguished from the optimal hopping distance as proposed by Mott’s formalism (discussed in later sections). Drift distance incidentally could be a cumulative result of several hops over large distances in the lattice. As seen in Fig. 2(c), the drift distance increases rapidly at small electric fields, and reaches a maximum as $E$ increases. Initially, $L_{\rm drift}(E)$ starts with a zero slope in the low disorder limit (blue, $W=0.2$) since localized states lie above the Fermi energy. In the range $0.040.02<E<$ 0.04 at $W=0.2$, the broadening of the band edge results in a rapid increase of the VRH transport. This is enabled by the tunneling of non-local reservoir electrons to localized states assisted by the electrostatic potential, similar to the Poole-Frenkel effect [43]. As the disorder increases, the threshold electric field required for the rapid growth of $L_{\rm drift}$ becomes smaller, reaching a critical condition at $W=0.4$ (red) - $0.45$ corresponding to the maximally VRH condition. As the disorder further increases ($W>0.5$), a significant fraction of sites are filled and the VRH mechanism gives way to strongly disordered transport with a much-reduced variation of $L_{\rm drift}$ by $E$, as discussed earlier. The crossover of the VRH transport to the Bloch oscillation behavior is most visible at the maximal VRH disorder at $W=0.4-0.45$, with the dashed line denoting the behavior of $x_{B}\propto 1/E$.
@@
Now, we turn our attention to more directly measurable quantities of conductance $\sigma(E)$ and mobility $\mu$. In Fig. 3(a), we show the conductivity averaged over 2000 disorder configurations from the finite-lattice calculations as a function of the electric field at different disorder strengths. At low disorder $W<0.4$ (red), the conductivity shows two distinct regimes. In the low-field regime, the constant conductivity is due to the transport of charge carriers that hybridize into the main lattice from the reservoirs. In the $W=0$ limit, the linear conductivity has an analytically exact expression [35]. As the field increases, the transport enters the activation regime (at $E\approx 0.015$ for $W=0$, black curve). As detailed in Appendix A, an analytic description (dashed line) of the activation behavior, Eq. (32), shows a reasonable agreement apart from an overall factor. As the disorder increases, the region of the linear conductivity shrinks until we get into the VRH regime close to $W=0.4\sim\Delta$. Once $W\gtrsim\Delta$, the conductivity increases very rapidly over many orders of magnitude up to $W\lesssim 2\Delta$, agreeing with the VRH range, Eq. (15). As the disorder increases further outside the VRH regime, the conductivity remains nearly independent of the field $E$, consistent with the Drude limit with a metallic weight as the system becomes delocalized.
@@
We gain a deeper understanding by studying the mobility $\mu$, i.e. the average conductivity per carrier density ($\mu=\sigma/\langle n_{l}\rangle$ as defined in Sec. II.1). The mobility, as shown in Fig. 3(b), maximizes in the VRH regime, defined in Eq. (15), in direct correspondence with the drift distance $L_{\rm drift}$ [see Fig. 2(d)]. As argued earlier, the enhanced mobility results from the transport through disorder-generated ingap states before the disorder potential becomes detrimental to the transport. The main difference of the mobility behavior from that of the drift distance [see Fig. 2(d)] is that the mobility decreases as $W\to 0$. This can be understood as follows. While the travel distance by the Bloch oscillation contributes to $L_{\rm drift}$, such oscillations cancel out in the DC transport for the conductivity and mobility. Therefore, mobility gives more relevant information for transport. The high disorder limit $\mu\sim W^{-2}$ (dashed line) is consistent with the Fermi-golden-rule interpretation.
@@
Figure 4: (a) The exponent $\nu$ is evaluated using nonlinear curve-fitting as described in the text. The computed exponent is very close to the theoretical value of $\nu=1/3$ depicted by the dashed black line for $W\lesssim W^{*}$. $\nu>1/3$ for $W>W^{*}$ since the disorder values fall outside of the VRH Regime. (b) Log of Conductivity is plotted against the electric field for different disorder strengths [$W\in\{0.2,0.3,0.35,0.4,0.45,0.5,0.6,0.7,0.8\}$: increasing in the order indicated by the arrow] with a fit (dashed line) of Eq. (20). The fit exponent is computed as $\nu$. (c) $\text{Log}(\sigma)$ plotted against $E^{-1/3}$ which depicts a straight line for all disorder values. We scale all the curves to match the fit of $W=0.4$ to shows a general agreement between the scaling law and the numerical result.
@@
We contrast the finite-lattice calculations with the CPA method. The results for conductivity in the CPA calculation are shown in Fig. 3(c). For weaker disorder strengths ($W<2\Delta=0.6$), the behavior is similar to the lattice calculations where the conductivity rises slowly at some value with respect to the electric field. For stronger disorder ($W>2\Delta$) the CPA-conductivity is strongly overestimated compared to the finite-lattice calculation, especially in the low-field limit. The insulator-to-metal transition occurs at $W=0.6$ for our chosen values of $\Delta$ and $\Gamma$ with the CPA conductivity decreasing with the field in the high disorder limit, in contrast to (a). As shown in Fig. 3(d), the mobility undergoes a sharp transition at $W=2\Delta$. This is the disorder strength at which the CPA spectral function band edge crosses the Fermi level. In contrast, the mobility from the finite-lattice calculations shows a gradual rise before capping at the critical disorder, where the Lifshitz tail smears into the Fermi sea at zero field limit.
@@
This contrast between both calculations indicates the difference in capturing the physics of the model. While the CPA method is generally successful in capturing the spectral properties at very high fields, it is unable to capture the transport physics of Anderson localization at lower electric fields hence it does not show any signatures of VRH which typically arise from the transport in localized levels.
III.2 Electric-Field Scaling Law of VRH
@@
In this section, we discuss an extension of Mott’s VRH behavior of Eq. (1) into the strong-field limit. The discussions on the strong-field scaling have been quite controversial. Differences in the $E$-field strength versus the temperature and the statistical criteria for hops under a field have led to competing predictions for the scaling exponent [28, 29, 32, 19]. Here, we present our own interpretation of the strong-$E$-field scaling and provide microscopic and unbiased lattice calculations to support our relation. The electric field accelerates the electrons which can be excited to non-local sites in a chain. We modify Mott’s model by incorporating the field-driven excitations in the Miller-Abraham’s hopping probability ${\cal W}$, Eq. (2), as
@@
with the hopping distance $x$ between two localized states of level-spacing $\epsilon$, the electric-field $E$, and the localization length $\xi$. The first term in the exponent captures the spatial overlap between two localized levels and the second term is modified to incorporate electric field excitations in place of the thermal excitation $k_{\text{B}}T$, as demonstrated in a dissipative field-driven lattice model [36]. At low temperatures, the electric field is the major source of excitations in the system and the scaling behavior may be dictated by the nonlinear electric field. Here, we imply that an electron gains the energy $eEx$ by traveling the distance $x$ in the direction of the field, and after making the hop, the excess energy relaxes into thermal energy. In the steady-state limit, repeated inelastic scattering establishes the effective temperature as a function of the most probable hopping distance $x$, which needs to be determined as below. As in the Drude model [44], we assume that energetic electrons relax their excess energy and thermalize before the next hop, often by emitting phonons. In our lattice calculations, the dissipation is provided by the hot-electron exchange with the fermion baths.
We minimize the function in the exponent of the Miller-Abraham’s relation, Eq. (16)
@@
In the strong-field limit $E\to\infty$, the electron motions are accelerated along the potential slope and the thermal factor can be ignored when the second term in $f(x)$ becomes small roughly when $Wa_{0}/(eEx^{2})\sim 1$ or $x\sim\sqrt{Wa_{0}/eE}\ll\xi$, beyond which the thermal factor becomes irrelevant. This threshold $x$ determines the most likely hopping distance and the Miller-Abraham’s formula takes the first term of the wave-function overlap [28, 29], and we approximate the conductivity as
@@
As in previous calculations, we perform straightforward numerical calculations on a finite-size electronic lattice with the electrostatic potential created by the field, by using the Keldysh Green’s function method. We then confirm the above prediction by performing a nonlinear curve-fit on our numerical simulation of the conductivity shown earlier in Fig. 3(a) to this scaling law in Eqs. (20) and (22). We compute the scaling exponent $\nu$ by treating $\sigma_{\infty}$, $E_{0}$ and $\nu$ as free parameters. The exponent $\nu$ is shown in Fig. 4(a) for different disorder strength. The estimated exponent is close to the theoretical value of $1/3$ predicted in Eq. (20) which is marked by the dashed line in Fig. 4(a). However, at strong disorders ($W>W^{*}$) where stronger electric fields are required for significant variations of conductivity, the numerically obtained exponent $\nu\approx 1/2$ becomes more consistent with Eq. (22).
@@
The nonlinear fit is depicted in Fig 4(b) which shows $\log({\sigma})$ against the electric field, with Eq. (20) shown as dashed lines. In the low-field limit, [$E<0.01$, this range is omitted in Fig. 4(b)], the dissipation effects dominate over the VRH transport and the conductivity is constant with electric field. While in the high-field limit ($E>0.4$) the fit overestimates the numerical conductivity as the system exhibits Bloch oscillations. Therefore, we note that Eq. (20) is a good fit at intermediate electric fields $0.13\lesssim E\lesssim 0.4$ where the VRH transport is observed. In Fig. 4(c), $\log({\sigma})$ vs $E^{-1/3}$ shows a linear behavior which captures the electric field scaling law in Eq. (20) reasonably well in the intermediate field regime.
@@
Figure 5: (a) Conductivity vs electric field ($\Gamma=0.0005$, $N=501$, $t=1$, $W=0.4$) in the VRH regime for different Temperatures. In the low-field limit, the constant part of the conductivity follows Mott’s $T^{-1/2}$-scaling of the VRH. Conductivity crosses over to a temperature-independent limit at high fields. The threshold field at which the conductance starts to deviate from the zero-field limit is evaluated using an empirical fit method detailed in the text, shown as the black dashed line for $T=0.02$. The threshold field progresses to higher values, as $T$ increases. (b) Threshold field $E_{\text{th}}$ vs temperature $T$.
We emphasize that the electric-field scaling is studied by performing quantum mechanical calculations on an electronic model without any reference to Mott’s statistical arguments. The derived scaling law applies to the regime in which the nonequilibrium excitations dominate the thermal effects of the bath, in contrast to statistical models [32, 33] and semi-classical approaches [27, 19]. The $E^{-1/3}$ scaling with electric field has been observed in some experiments [46, 47] while the exponent $\nu=1/2$ also has been reported [30, 48]. In the literature, different theoretical assumptions have led to different power-law scaling behaviors [33, 49, 48, 20, 50, 51], and our calculations directly based on quantum mechanical equations of motion may offer some insights into the nonlinear VRH transport in low dimensional disordered systems.
@@
We conclude this section by presenting experimentally relevant temperature dependence of the VRH in the high-field regime. A log-log plot of the conductivity vs the electric field, shown in Fig. 5(a), displays two transport regime separated by the crossover behavior. The low-field behavior of the constant conductivity is governed by the linear conductivity, which is consistent with Mott’s $T^{-1/2}$-scaling. The high-field limit, however, is nearly independent of the temperature. This is again consistent with the view that a high $E$ destroys the localization. Constrained by the two limits, the temperature dependence of the localization-delocalization crossover is of a strong experimental interest. The debate also relates to resistive switching, the case of more distinct insulator-to-metal transitions driven by an electric field, where the relation between the threshold field and the temperature may point to the nature of the transition as classical vs quantum [52].
@@
Our scenario based on the disorder model suggests something less dramatic. The $\sigma_{1}(T)$-term strongly varies with temperature over a few orders of magnitude due to the VRH mechanism, while the high-$E$ limit of the conductivity $\sigma(E,T)$ has much weaker $T$-dependence since the noneqiulibrium-driven effective temperature would overwhelm the bath temperature. Then, by writing the conductivity as $\sigma_{1}(T)+\sigma^{\prime}(E)$ where $\sigma^{\prime}(E)$ is $T$-independent and an increasing function of $E$, the crossover would happen when $\sigma_{1}(T)\approx\sigma^{\prime}(E_{\rm th})$, thus $E_{\rm th}$ is an increasing function of $T$.

### v3

@@ III Results
III.1 Electron Drift in Disordered Lattice
@@
We compute $C(x)$ in the disordered-lattice calculation numerically by computing the disorder-averaged correlation of the non-diagonal matrix elements terms of the retarded Green’s function at the local Fermi energy mE-mE averaged over each $m$-th site, since the transport is carried over states close to the Fermi energy. Here, $x$ denotes the spatial index ($0\leq x\leq 150$) across the 1D lattice and $M=201$ denotes the number of sites chosen for spatial averaging around the center of the lattice of $N=501$ sites. This quantity $C(x)$ is averaged over multiple disorder configurations ($\sim 2000$ different configurations).
For strongly localized states, this quantity decays exponentially and for a delocalized or extended state, it spreads over the entire lattice. Fig. 2(a) depicts the behavior of $C(x)$ for increasing electric field at the disorder strength $W=0.4$ inside the VRH regime. At low electric fields, the wave functions are typically localized; but at higher fields, the wave function shows evidence for the Bloch oscillations. Although the Bloch oscillation localizes the electron motion via the Brillouin zone averaging, it is only possible when the electron traverses the Brillouin zone. Therefore, the emergence of the Bloch oscillation can be taken as a sign of delocalization. The spread of $C(x)$ behaves inversely proportional to the electric field and at high fields $E\gtrsim 0.4$, $C(x)$ extends to the distance $x_{B}=4t/E$ (as marked by an arrow for $E=0.07$). The width of the Bloch oscillation can be easily understood as the travel distance of an electron driven by an electric field as
@@
with the group velocity $v_{g}(k)=2t\sin(k)$ and the semi-classical equation of motion $\hbar\dot{k}=eEa$. This behavior corroborates with the discussion [34] that, at high fields, the agreement of the CPA and the disordered-lattice spectral functions indicates the tendency of delocalization.
@@
Figure 3: (a) Conductivity vs electric field ($\Gamma=0.0005$, $N=501$, $t=1$, $T=0.01$). The average conductivity slowly increases with the electric field. In the low electric field region the conductivity strongly varies with the disorder strengths $W\in\{0.2,0.3,0.35,0.4,0.45,0.5,0.6,0.7,0.8,1.0,1.2\}$: increasing in order indicated by the arrow in 3(b)] with a rapid rise in the VRH regime $0.3<W<0.5$. The black solid and the dashed line show the numerical and analytic behaviors [see Eq. (34)], respectively, at zero disorder. (b) Mobility vs disorder strength ($\Gamma=0.0005$, $N=501$, $t=1$, $T_{B}=0.01$) for different electric fields. The average mobility at low electric fields $E<0.04$ sharply rises in the disorder range $\Delta<W\lesssim 0.5$ marked by the black dashed lines for the VRH regime. Beyond this disorder range ($W>0.5$) the mobility typically follows roughly the Fermi golden-rule ($\sim W^{-2}$) shown by the black dashed line. (c) CPA calculation of conductivity. While the overall behavior is consistent with the disordered-lattice calculation, the high-disorder limit strongly overestimates the conductivity. (d) CPA calculation of mobility. Despite the similarity of the overall lineshape, the CPA overestimates the mobility and fails to produce the correct trends in the high-$W$ limit.
@@
$L_{\rm drift}=\frac{\int_{0}^{\infty}C(x)xdx}{\int_{0}^{\infty}C(x)dx}.$ (16)
We make a note here that the drift distance is the travel distance of wave-propagation under the effect of an external field during the charge carrier’s lifetime. This is distinguished from the optimal hopping distance as proposed by Mott’s formalism (discussed in later sections). Drift distance incidentally, could be a cumulative result of several hops over large distances in the lattice. As seen in Fig. 2(c), the drift distance increases rapidly at small electric fields, and reaches a maximum as $E$ increases. Initially, $L_{\rm drift}(E)$ starts with a zero slope in the low disorder limit (blue, $W=0.2$) since localized states lie above the Fermi energy. In the range $0.02<E<0.04$ at $W=0.2$, the broadening of the band edge results in a rapid increase of the VRH transport. This is enabled by the tunneling of non-local reservoir electrons to localized states assisted by the electrostatic potential, similar to the Poole-Frenkel effect [42]. As the disorder increases, the threshold electric field required for the rapid growth of $L_{\rm drift}$ becomes smaller, reaching a critical condition at $W=0.4$ (red) - $0.45$ corresponding to the maximally VRH condition. As the disorder further increases ($W>0.5$), a significant fraction of sites are filled and the VRH mechanism gives way to strongly disordered transport with a much-reduced variation of $L_{\rm drift}$ by $E$, as discussed earlier. The crossover of the VRH transport to the Bloch oscillation behavior is most visible at the maximal VRH disorder at W=0.40.45W=0.4-0.45, with the dashed line denoting the behavior of $x_{B}\propto 1/E$.
@@
III.2 Nonlinear Conductivity in Disordered Lattice
Now, we turn our attention to more directly measurable quantities of conductivity $\sigma(E)$ and mobility $\mu$. In Fig. 3(a), we show the conductivity averaged over 2000 disorder configurations from the disordered-lattice calculations as a function of the electric field at different disorder strengths. At low disorder $W<0.4$ (red), the conductivity shows two distinct regimes, separated by the shoulders in the range $E=0.01\sim 0.02$. In the high-field regime, the conductivity shows the activation behavior by the electric field, mainly given by the Landau-Zener tunneling [43]. The exponential behavior of the conductivity is overridden by the coupling to the reservoirs, and it reaches a plateau as $E\to 0$. As discussed in previous work [34], the conduction electrons hybridize with the reservoir continuum for the electron density $n\propto\Gamma$ with the acceleration time proportional to $\Gamma/\Delta$, leading to $\sigma(E\to 0)\propto\Gamma^{2}$ [Eq. (19) of Ref.[34]]. Despite the particular form of dissipation in our model, the coupling to the environment is universal, and the lower-bound behavior of the field-activation is generally expected.
@@ III.2 Nonlinear Conductivity in Disordered Lattice
As detailed in Appendix-A, an analytic description (dashed line) of the activation behavior, Eq. (34), shows a reasonable agreement apart from an overall factor. As the disorder increases, the region of the linear conductivity shrinks until we get into the VRH regime close to $W=0.4\sim\Delta$. Once $W\gtrsim\Delta$, the conductivity increases very rapidly over many orders of magnitude up to $W\lesssim 2\Delta$, agreeing with the VRH range, Eq. (17). As the disorder increases further outside the VRH regime, the conductivity remains nearly independent of the field $E$, consistent with the Drude limit with a metallic weight as the system becomes delocalized.
@@
We gain a deeper understanding by studying the mobility $\mu$, i.e. the average conductivity per carrier density ($\mu=\sigma/\langle n_{l}\rangle$ as defined in section-II.1). The mobility, as shown in Fig. 3(b), maximizes in the VRH regime, defined in Eq. (17), in direct correspondence with the drift distance $L_{\rm drift}$ [see Fig. 2(d)]. As argued earlier, the enhanced mobility results from the transport through disorder-generated ingap states before the disorder potential becomes detrimental to the transport. The main difference of the mobility behavior from that of the drift distance [see Fig. 2(d)] is that the mobility decreases as $W\to 0$. This can be understood as follows. While the travel distance by the Bloch oscillation contributes to $L_{\rm drift}$, such oscillations cancel out in the DC transport for the conductivity and mobility. Therefore, mobility gives more relevant information for disordered transport. The high disorder limit $\mu\sim W^{-2}$ (dashed line) is consistent with the Fermi-golden-rule interpretation.
@@
Figure 4: (a) The exponent $\nu$ is evaluated using nonlinear curve-fitting as described in the text. The computed exponent is very close to the theoretical value of $\nu=1/3$ depicted by the dashed black line for $W\lesssim W^{*}$. $\nu>1/3$ for $W>W^{*}$ since the disorder values fall outside of the VRH Regime. (b) Log of Conductivity is plotted against the electric field for different disorder strengths $W\in\{0.2,0.3,0.35,0.4,0.45,0.5,0.6,0.7,0.8\}$: increasing in the order indicated by the arrow] with a fit (dashed line) of Eq. (22). The fit exponent is computed as $\nu$. (c) $\text{Log}(\sigma)$ plotted against $E^{-1/3}$ which depicts a straight line for all disorder values. We scale all the curves to match the fit of $W=0.4$ to show a general agreement between the scaling law and the numerical result.
@@
We contrast the disordered-lattice calculations with the CPA method. The results for conductivity in the CPA calculation are shown in Fig. 3(c). For weaker disorder strengths ($W<2\Delta=0.6$), the behavior is similar to the lattice calculations, where the conductivity rises slowly at some value with respect to the electric field. For stronger disorder ($W>2\Delta$), the CPA-conductivity is strongly overestimated compared to the disordered-lattice calculation, especially in the low-field limit. The insulator-to-metal transition occurs at $W=0.6$ for our chosen values of $\Delta$ and $\Gamma$, with the CPA conductivity decreasing with the field in the high disorder limit, in contrast to (a). As shown in Fig. 3(d), the mobility undergoes a sharp transition at $W=2\Delta$. This is the disorder strength at which the CPA spectral function band edge crosses the Fermi level[34]. In contrast, the mobility from the disordered-lattice calculations shows a gradual rise before capping at the critical disorder, where the Lifshitz tail smears into the Fermi sea at the zero-field limit.
@@
This contrast between the two calculations indicates the difference in capturing the physics of the model. While the CPA method is generally successful in capturing the spectral properties at very high fields, it is unable to capture the transport physics of Anderson localization at lower electric fields hence, it does not show any signatures of VRH, which typically arise from the transport in localized levels.
III.3 Electric-Field Scaling Law of VRH
@@
In this section, we discuss an extension of Mott’s VRH behavior of Eq. (1) into the strong-field limit. The discussions on the strong-field scaling have been quite controversial. Differences in the $E$-field strength versus the temperature and the statistical criteria for hops under a field have led to competing predictions for the scaling exponent [27, 28, 31, 19]. Here, we present our own interpretation of the strong-$E$-field scaling and provide microscopic and unbiased lattice calculations to support our relation. The electric field accelerates the electrons, which can be excited to non-local sites in a chain. We modify Mott’s model by incorporating the field-driven excitations in the Miller-Abraham’s hopping probability $\cal W$}, Eq. (2), as
@@
with the hopping distance $x$ between two localized states of level-spacing $\epsilon$, the electric-field $E$, and the localization length $\xi$. The first term in the exponent captures the spatial overlap between two localized levels, and the second term is modified to incorporate electric field excitations in place of the thermal excitation $k_{\text{B}}T$, as demonstrated in a dissipative field-driven lattice model [35]. At low temperatures, the electric field is the major source of excitations in the system, and the scaling behavior may be dictated by the nonlinear electric field. Here, we imply that an electron gains the energy $eEx$ for a most probable hopping distance $x$ in the direction of the field, and after making the hop, the excess energy relaxes into thermal energy. The driven chain thermalizes by emission of hot-electrons followed by absorption of cold-electrons from the reservoirs, and by creating energetic electron-hole pairs in the infinite sea of electrons in the Fermion bath. It has been shown analytically [43] that the electron’s effective temperature is linearly proportional to the field in the clean limit in the particle-hole symmetric metals and insulators. Our single-band insulator model, which lacks the particle-hole symmetry, shows a similar trend of increasing electron temperature, as observed from the smearing of the nonequilibrium distribution with increasing electric field in our previous study [Fig. 5(a) of [34]].
In the steady-state limit, repeated inelastic scattering establishes an effective temperature as a function of the most probable hopping distance $x$. With this thermal description of the field excitation, we proceed analogously with Mott’s mechanism of VRH, for the $E$-field scaling relation. We minimize the function in the exponent of the Miller-Abraham’s relation, Eq. (18)
@@
Discussions so far apply in the intermediate field regime where the balance between the two terms in Eq. (21) leads the probable hopping range $x^{*}$ to be comparable to $\xi$, or equivalently, $T_{\rm eff}=eE\xi/3\sim Wa_{0}/(3\xi)$ where the electron excitation energy by the field remains a fraction of the disorder potential strength $W$.
In a stronger field limit $E\to\infty$, however, the field becomes so strong that the potential drop by the field immediately makes most hops energetically favorable. This regime is reached when $Wa_{0}/(eEx^{2})\sim 1$ or $x\sim\sqrt{Wa_{0}/eE}\ll\xi$, beyond which the thermal factor becomes irrelevant. This threshold $x$ determines the most likely hopping distance, and the Miller-Abraham’s formula takes the first term of the wave-function overlap [27, 28], and we approximate the conductivity as
@@
Now, we present the numerical results. As in previous calculations, we perform straightforward numerical calculations on a finite-size electronic lattice with the electrostatic potential created by the field, by using the Keldysh Green’s function method. We then confirm the above prediction by performing a nonlinear curve-fit on our numerical simulation of the conductivity shown earlier in Fig. 3(a) to this scaling law in Eqs. (22) and (24). We compute the scaling exponent $\nu$ by treating $\sigma_{\infty}$, $E_{0}$ and $\nu$ as free parameters. The exponent $\nu$ is shown in Fig. 4(a) for different disorder strengths. The estimated exponent is close to the theoretical value of $1/3$ predicted in Eq. (22), which is marked by the dashed line in Fig. 4(a). However, at strong disorders ($W>W^{*}$) where stronger electric fields are required for significant variations of conductivity, the numerically obtained exponent $\nu\approx 1/2$ becomes more consistent with Eq. (24).
@@
The nonlinear fit is depicted in Fig. 4(b), which shows $\log({\sigma})$ against the electric field, with Eq. (22) shown as dashed lines. In the low-field limit, [$E<0.01$, this range is omitted in Fig. 4(b)], the dissipation effects dominate over the VRH transport, and the conductivity is constant with electric field. While in the high-field limit ($E>0.4$), the fit overestimates the numerical conductivity as the system exhibits Bloch oscillations. Therefore, we note that Eq. (22) is a good fit at intermediate electric fields $0.13\lesssim E\lesssim 0.4$ where the VRH transport is observed. In Fig. 4(c), $\log({\sigma})$ vs $E^{-1/3}$ shows a linear behavior which captures the electric field scaling law in Eq. (22) reasonably well in the intermediate field regime.
@@
Figure 5: (a) Conductivity vs electric field ($\Gamma=0.0005$, $N=501$, $t=1$, $W=0.4$) in the VRH regime for different Temperatures. In the low-field limit, the constant part of the conductivity follows Mott’s $T^{-1/2}$-scaling of the VRH. Conductivity crosses over to a temperature-independent limit at high fields. The threshold field at which the conductance starts to deviate from the zero-field limit is evaluated using an empirical fit method detailed in the text, shown as the black dashed line for $T=0.02$. The threshold field progresses to higher values as $T$ increases. (b) Threshold field $E_{\text{th}}$ vs temperature $T$.
Different experiments have reported different scaling exponents [32, 45, 46, 20, 47, 48], with the $E^{-1/3}$ scaling been observed in some experiments [49, 50] and $E^{-1/2}$ scaling most commonly reported [46, 45, 47]. Although various theoretical assumptions[31, 32, 26, 19, 29] have been used to explain these scaling behaviors, based on material properties under study. However, the microscopic mechanism of variable range hopping under an electric field is still a matter of debate. Our calculations, apart from the choice of the model, are directly based on quantum mechanical equations of motion and offer insights into the nonlinear VRH transport in such low-dimensional disordered systems. The electronic model of a one-dimensional disordered chain presented in this study could be reflected in some low-dimensional disordered semiconductors [51], which have been used to study nonlinear transport under a varied range of electric fields.
@@
We conclude this section by presenting experimentally relevant temperature dependence of the VRH in the high-field regime. The temperature dependence of the threshold behavior is often considered an important clue to various resistive transitions for the debate of electronic versus thermal mechanism [52, 53]. A log-log plot of the conductivity vs the electric field, shown in Fig. 5(a), displays two transport regimes separated by the crossover behavior. The low-field behavior of the constant conductivity is governed by the linear conductivity, which is consistent with Mott’s $T^{-1/2}$-scaling. The high-field limit, however, is nearly independent of the temperature. This is again consistent with the view that a high $E$ destroys the localization. Constrained by the two limits, the temperature dependence of the localization-delocalization crossover is of strong experimental interest.
@@
Our scenario, based on the disorder model, suggests something less dramatic. The $\sigma_{1}(T)$-term strongly varies with temperature over a few orders of magnitude due to the VRH mechanism, while the high-$E$ limit of the conductivity $\sigma(E,T)$ has much weaker $T$-dependence since the nonequilibrium-driven effective temperature would overwhelm the bath temperature. Then, by writing the conductivity as $\sigma_{1}(T)+\sigma^{\prime}(E)$ where $\sigma^{\prime}(E)$ is $T$-independent and an increasing function of $E$, the crossover would happen when $\sigma_{1}(T)\approx\sigma^{\prime}(E_{\rm th})$, thus $E_{\rm th}$ is an increasing function of $T$.